OpenStudy (anonymous):

3. What is f(–3) if f(x) = |2x + 1| + 5? (1 point) 0 10 12 –10

OpenStudy (anonymous):

To find f(-3) just plug in x=-3. You will get f(-3)=10.Remember that mort of something always gives the absolute value of that something. You can check for yourself. Hope this helps you.

OpenStudy (anonymous):

Thank you :)

OpenStudy (kc_kennylau):

\[\begin{array}{cl} &f(-3)\\ =&|2(-3)+1|+5\\ =&|-6+1|+5\\ =&|-5|+5\\ =&5+5\\ =&10 \end{array}\]